### v2

# On an unconditional $\mathrm{GL}_{3}$ analog of Selberg’s result

Abstract.
Let $F$ be a Hecke–Maass cusp form for $\mathrm{SL}_{3}(\mathbb{Z})$ with the Langlands parameter $\mu_{F}=\big(\mu_{F,1},\mu_{F,2},\mu_{F,3}\big)$ and the associated $L$-function $L(s,F)$. Define $S_{F}(t)=\pi^{-1}\arg L(1/2+\mathrm{i}t,F)$. When $\mu_{F}$ is in generic position, we establish an unconditional asymptotic formula for the moments of $S_{F}(t)$. Previously, such a formula was only known to hold under the Generalized Riemann Hypothesis. The key ingredient is a weighted zero-density estimate in the spectral aspect for $L(s,F)$, which has recently been proved by Sun and Wang in arXiv:2412.02416.
Key words and phrases:
Moments, $\mathrm{GL}_{3}$ Hecke–Maass cusp forms, Selberg’s limit theorem.
2010 Mathematics Subject Classification:
11F12, 11F66, 11F72.
†  Corresponding author
1. Introduction
Let $L(s,F)$ be an $L$-function and define
$\begin{split}S_{F}(t)=\frac{1}{\pi}\arg L\Big(\frac{1}{2}+\mathrm{i}t,F\Big).\end{split}$
Here, the argument is defined by continuous variation along the horizontal segment from $\infty+\mathrm{i}t$ to $1/2+\mathrm{i}t$, commencing at $\infty+\mathrm{i}t$ with the value 0. It is very interesting to study the function $S_{F}(t)$ in number theory. For example, in the case of the Riemann zeta function, the classical Riemann–von Mangoldt formula states that the number of zeros $\rho=\beta+\mathrm{i}\gamma$ of $\zeta(s)$ (counted with multiplicity) in the rectangle $\{(\beta,\gamma)\mid 0<\beta<1,0<\gamma\leq t\}$ is given by
$\frac{t}{2\pi}\log\frac{t}{2\pi\mathrm{e}}+\frac{7}{8}+\frac{1}{\pi}\arg\zeta(1/2+\mathrm{i}t)+O\Big(\frac{1}{t}\Big),$
which implies that the function $S_{\zeta}(t)=\pi^{-1}\arg\zeta(1/2+\mathrm{i}t)$ is closely related to the nontrivial zeros of $\zeta(s)$. As such, the fascinating behavior of $S_{\zeta}(t)$ has been intensively studied by many number theorists (see, e.g., [1, 24, 5, 8] and the references therein). On the other hand, the function $S_{F}(t)$ has a close connection with random matrix theory (see [2]). In fact, the behavior of $S_{F}(t)$ is analogous to the extreme value statistics of the characteristic polynomials of random matrices.
To understand more about $S_{F}(t)$, an effective way is to calculate and analyze the various moments of $S_{F}(t)$. In the 1940s, Selberg [21, 22] examined the distribution of $S_{\zeta}(t)$ and demonstrated that for any $n\in\mathbb{N}$,
$\frac{1}{T}\int_{T}^{2T}S_{\zeta}(t)^{2n}\mathrm{d}t=\frac{(2n)!}{n!(2\pi)^{2n}}(\log\log T)^{n}+O\big((\log\log T)^{n-1/2}\big),$ (1.1)
which implies that, in an average sense, $|S_{\zeta}(t)|$ has an order of magnitude $\sqrt{\log\log T}$. Additionally, for the Dirichlet $L$-function $L(s,\chi)$ with primitive $\chi$, Selberg [23] studied the behavior of $S_{\chi}(t)=\frac{1}{\pi}\arg L\big(\frac{1}{2}+\mathrm{i}t,\chi\big)$ and obtained, for prime modulus $q$ and any fixed $t>0$,
$\frac{1}{q}\,\,\sideset{}{{}^{\star}}{\sum}_{\chi\bmod q}S_{\chi}(t)^{2n}=\frac{(2n)!}{n!(2\pi)^{2n}}(\log\log q)^{n}+O\big((\log\log q)^{n-1/2}\big),$ (1.2)
where $\star$ indicates that the summation runs over the primitive characters $\chi$ modulo $q$. It should be noted that the counterparts of (1.1) and (1.2) for the odd-order moments can also be derived by making minor adjustments to Selberg’s approach.
For the $\mathrm{GL}_{2}$ family of $L$-functions, Hejhal and Luo [11] first considered the generalization of (1.2) to the case of eigenfunctions $\psi$ of the hyperbolic Laplacian $\Delta$ for the full modular group. More precisely, let $L(s,\psi)$ be the standard Hecke–Maass $L$-function associated to $\psi$ and define
$S_{\psi}(t)=\frac{1}{\pi}\arg L\Big(\frac{1}{2}+\mathrm{i}t,\psi\Big).$
They established the analogue of (1.1) when $\psi$ is an Eisenstein series. In addition, when $\psi$ is the Hecke–Maass cusp form $u_{j}$ of $\mathrm{SL}_{2}(\mathbb{Z})$ with the first Fourier coefficient $\nu_{j}(1)$ and Laplacian eigenvalue $\lambda_{j}=1/4+t_{j}^{2}$ ($t_{j}\geq 0$), under the Generalized Riemann Hypothesis (GRH), they showed that for any fixed $t,\beta>0$ and any large positive parameter $T$,
$\begin{split}\sum_{j=1}^{\infty}\mathrm{e}^{-\beta(t_{j}-T)^{2}}|\nu_{j}(1)|^{2}S_{u_{j}}(t)^{2n}=\frac{2\pi^{-3/2}}{\sqrt{\beta}}\frac{(2n)!}{n!(2\pi)^{2n}}T(\log\log T)^{n}+O_{t,n,\beta}\big(T(\log\log T)^{n-1/2}\big).\end{split}$
Recently, Sun and Wang [19] reestablished this asymptotic formula by means of a weighted zero-density estimate instead of assuming the GRH. Moreover, for the holomorphic case, Liu and Shim [16] and Sun et al. [18] successively obtained unconditional analogues in the weight aspect and in the level aspect, respectively.
For the $\mathrm{GL}_{3}$ family of $L$-functions, Liu and Liu [15] explored the average behavior of the analogue function $S_{F}(t)$, where $F$ is a Hecke–Maass cusp form for $\mathrm{SL}_{3}(\mathbb{Z})$. To state their results, we need to introduce some notations. For a cusp form $F$ in an orthogonal basis of Hecke–Maass cusp forms for $\mathrm{SL}_{3}(\mathbb{Z})$, we denote its Langlands parameter by $\mu_{F}=\big(\mu_{F,1},\mu_{F,2},\mu_{F,3}\big)$, where $\mu_{F}$ is a point in the region
$\begin{split}\Lambda_{1/2}^{\prime}=\Bigg\{(\mu_{F,1},\mu_{F,2},\mu_{F,3})\in\mathbb{C}^{3},\,\,\,\begin{aligned} &\mu_{F,1}+\mu_{F,2}+\mu_{F,3}=0,\,\,|\operatorname{Re}(\mu_{F,j})|\leq\frac{1}{2},\,j=1,2,3,\\ &\{-\mu_{F,1},-\mu_{F,2},-\mu_{F,3}\}=\{\overline{\mu_{F,1}},\overline{\mu_{F,2}},\overline{\mu_{F,3}}\}\end{aligned}\Bigg\}\end{split}$
in the Lie algebra $\mathfrak{sl}_{3}(\mathbb{C})$. Denote by $\nu_{F}=\big(\nu_{F,1},\nu_{F,2},\nu_{F,3}\big)$ the spectral parameter of $F$, where
$\begin{split}\nu_{F,1}:=\frac{1}{3}&\big(\mu_{F,1}-\mu_{F,2}\big),\quad\nu_{F,2}:=\frac{1}{3}\big(\mu_{F,2}-\mu_{F,3}\big),\\ \nu_{F,3}:=&-\nu_{F,1}-\nu_{F,2}=\frac{1}{3}\big(\mu_{F,3}-\mu_{F,1}\big).\end{split}$ (1.3)
Then we have
$\mu_{F,1}=2\nu_{F,1}+\nu_{F,2},\quad\mu_{F,2}=\nu_{F,2}-\nu_{F,1},\quad\mu_{F,3}=-\nu_{F,1}-2\nu_{F,2}.$ (1.4)
We will use the coordinates $\mu$ and $\nu$ simultaneously. Note that $\big(\nu_{F,1}-1/3,\nu_{F,2}-1/3\big)$ is the type of $F$ as defined in Goldfeld’s book [10]. Assume that $F$ is normalized so that the $(1,1)$-th Fourier coefficient $A_{F}(1,1)=1$. The $L$-function associated to $F$ is defined by
$L(s,F)=\sum_{n=1}^{\infty}\frac{A_{F}(1,n)}{n^{s}},\quad\,\operatorname{Re}(s)>1,$
which has the Euler product
$\begin{split}L(s,F)&=\prod_{p}\bigg(1-\frac{A_{F}(1,p)}{p^{s}}+\frac{A_{F}(p,1)}{p^{2s}}-\frac{1}{p^{3s}}\bigg)^{-1}\\ &=\prod_{p}\big(1-\alpha_{F,1}(p)p^{-s}\big)^{-1}\big(1-\alpha_{F,2}(p)p^{-s}\big)^{-1}\big(1-\alpha_{F,3}(p)p^{-s}\big)^{-1}.\end{split}$
Here we have
$\begin{split}\begin{cases}\alpha_{F,1}(p)\alpha_{F,2}(p)\alpha_{F,3}(p)=1,\\ \alpha_{F,1}(p)+\alpha_{F,2}(p)+\alpha_{F,3}(p)=A_{F}(1,p),\\ \alpha_{F,1}(p)\alpha_{F,2}(p)+\alpha_{F,2}(p)\alpha_{F,3}(p)+\alpha_{F,3}(p)\alpha_{F,1}(p)=A_{F}(p,1),\end{cases}\end{split}$ (1.5)
and by Kim and Sarnak [13],
$\begin{split}|\alpha_{F,j}(p)|\leq p^{\frac{5}{14}},\quad\text{and}\quad|\operatorname{Re}(\mu_{F,j})|\leq\frac{5}{14},\quad\text{for}\,\,j=1,2,3.\end{split}$ (1.6)
Note that $L(s,F)$ has an analytic continuation to the whole complex plane and satisfies the functional equation
[…]
where $\Gamma_{\mathbb{R}}(s)=\pi^{-s/2}\Gamma(s/2)$ and the dual form of $F$ is denoted by $\widetilde{F}$ with the Langlands parameter $(-\mu_{F,1},-\mu_{F,2},-\mu_{F,3})$, the spectral parameter $\big(\nu_{F,2},\nu_{F,1},\nu_{F,3}\big)$ and the coefficients $A_{\widetilde{F}}(m,n)=\overline{A_{F}(m,n)}=A_{F}(n,m)$.
Let $\mu_{0}=(\mu_{0,1},\mu_{0,2},\mu_{0,3})$ and $\nu_{0}=(\nu_{0,1},\nu_{0,2},\nu_{0,3})$, which satisfy the corresponding relations (1.3) and (1.4). We also assume that $\mu_{0}$ is in generic position, i.e.,
$|\mu_{0,j}|\asymp|\nu_{0,j}|\asymp T:=\|\mu_{0}\|,\quad 1\leq j\leq 3,$ (1.8)
where
$\|\mu_{0}\|=\sqrt{|\mu_{0,1}|^{2}+|\mu_{0,2}|^{2}+|\mu_{0,3}|^{2}}.$
Throughout this paper, we let $M\!=\!T^{1-\eta}$ for some sufficiently small $0<\eta<1/2$. Define a test function $h_{T,M}(\mu)$ (depending on $\mu_{0}$) for $\mu=(\mu_{1},\mu_{2},\mu_{3})\in\mathbb{C}^{3}$ by
$h_{T,M}(\mu):=P(\mu)^{2}\bigg(\sum_{w\in\mathcal{W}}\psi\bigg(\frac{w(\mu)-\mu_{0}}{M}\bigg)\bigg)^{2},$
where $\psi(\mu)=\exp\big(\mu_{1}^{2}+\mu_{2}^{2}+\mu_{3}^{2}\big)$ and
$P(\mu)=\prod_{0\leq n\leq A}\prod_{j=1}^{3}\frac{\big(\nu_{j}-\frac{1}{3}(1+2n)\big)\big(\nu_{j}+\frac{1}{3}(1+2n)\big)}{|\nu_{0,j}|^{2}}$
for some fixed large $A>0$. Here
$\mathcal{W}:=\Big\{I,\;w_{2}=\Big(\begin{smallmatrix}1&&\\ &&-1\\ &1&\end{smallmatrix}\Big),\;w_{3}=\Big(\begin{smallmatrix}&1&\\ -1&&\\ &&1\end{smallmatrix}\Big),\;w_{4}=\Big(\begin{smallmatrix}&1&\\ &&1\\ 1&&\end{smallmatrix}\Big),\;w_{5}=\Big(\begin{smallmatrix}&&1\\ 1&&\\ &1&\end{smallmatrix}\Big),\;w_{6}=\Big(\begin{smallmatrix}&&1\\ &-1&\\ 1&&\end{smallmatrix}\Big)\Big\}$
is the Weyl group of $\mathrm{SL}_{3}(\mathbb{R})$. The function $h_{T,M}(\mu)$ localizes at a ball of radius $M$ about $w(\mu_{0})$ for each $w\in\mathcal{W}$.
Moreover, we define the spectral measure on the hyperplane $\mu_{1}+\mu_{2}+\mu_{3}=0$ by
$\mathrm{d}_{\text{spec}}\mu=\text{spec}(\mu)\mathrm{d}\mu,$
where
$\text{spec}(\mu):=\prod_{j=1}^{3}\bigg(3\nu_{j}\tan\bigg(\frac{3\pi}{2}\nu_{j}\bigg)\bigg)\quad\mbox{and}\quad\mathrm{d}\mu=\mathrm{d}\mu_{1}\mathrm{d}\mu_{2}=\mathrm{d}\mu_{2}\mathrm{d}\mu_{3}=\mathrm{d}\mu_{1}\mathrm{d}\mu_{3}.$
By a trivial estimate, we have
$\mathcal{H}:=\frac{1}{192\pi^{5}}\int_{\operatorname{Re}(\mu)=0}h_{T,M}(\mu)\mathrm{d}_{\rm{spec}}\mu\asymp T^{3}M^{2}.$ (1.9)
Fix any $t>0$ and $n\in\mathbb{N}$, under the GRH, Liu and Liu [15] established that, for large positive $T$,
$\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\nu_{F})}{L(1,\text{Ad}^{2}F)}S_{F}(t)^{2n}=\frac{9(2n)!}{4n!(2\pi)^{2n}}(\log\log T)^{n}+O_{t,n}\big((\log\log T)^{n-1/2}\big),\end{split}$ (1.10)
where $L(s,\text{Ad}^{2}F)$ denotes the adjoint square $L$-function of $F$. Recently, with the automorphic Plancherel density theorem of Matz and Templier, Chen et al. [9] extended their result to $\mathrm{GL}_{n}$. In this paper, we are devoted to proving the asymptotic formula (1.10) without assuming the GRH. We introduce the normalizing factor
$\mathcal{N}_{F}:=\|F\|^{2}\prod_{j=1}^{3}\cos\bigg(\frac{3}{2}\pi\nu_{F,j}\bigg).$
By Rankin–Selberg theory in combination with Stade’s formula (see e.g., Blomer [4, Section 4] and Li [14, Theorem 2]), it is easy to see that
$\begin{split}\mathcal{N}_{F}\asymp\mathrm{Res}_{s=1}L(s,F\times\widetilde{F})\ll\|\mu_{F}\|^{\varepsilon},\end{split}$
[…]
$\begin{split}L(s,F\times\widetilde{F})&=\zeta(3s)\sum_{m,n\geq 1}\frac{|A_{F}(m,n)|^{2}}{(m^{2}n)^{s}}\\ &=\prod_{p}\prod_{i=1}^{3}\prod_{j=1}^{3}\bigg(1-\alpha_{F,i}(p)\overline{\alpha_{F,j}(p)}p^{-s}\bigg)^{-1},\quad\text{for}\,\,\,\operatorname{Re}(s)>1,\end{split}$
and the implied constants depend at most on $\varepsilon$. Our main result is the following theorem.
Theorem 1.1.
Let $t>0$ be fixed and $T$ be any large positive parameter. We have
$\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}S_{F}(t)^{n}=C_{n}(\log\log T)^{n/2}+O_{t,n}\big((\log\log T)^{(n-1)/2}\big),\end{split}$ (1.11)
where
$C_{n}=\begin{cases}\frac{n!}{(n/2)!(2\pi)^{n}},\,&\textit{if }\,n\,\,\text{is even},\\ 0,\,&\textit{if }\,n\,\,\text{is odd}.\end{cases}$
Moreover, define the following probability measure $\mu_{TM}$ on $\mathbb{R}$ by
$\begin{split}\mu_{TM}(E):=\bigg(\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\mathbf{1}_{E}\bigg(\frac{S_{F}(t)}{\sqrt{\log\log T}}\bigg)\bigg)\bigg/\bigg(\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg),\end{split}$
where $\mathbf{1}_{E}$ is the characteristic function of a Borel measurable set $E$ in $\mathbb{R}$. Then for every $a<b$, we have
$\lim_{T\rightarrow\infty}\mu_{TM}([a,b])=\sqrt{\pi}\int_{a}^{b}\exp(-\pi^{2}\xi^{2})\mathrm{d}\xi,$
uniformly with respect to $M$.
Remark 1.
For $F$ a Hecke–Maass cusp form of $\mathrm{SL}_{3}(\mathbb{Z})$, our result indicates that, as $T\rightarrow\infty$, the ratios $S_{F}(t)/\sqrt{\log\log T}$ converge in distribution to a normal distribution of mean 0 and variance $(2\pi^{2})^{-1}$.
Notation Throughout the paper, $\varepsilon$ is an arbitrarily small positive real number, which may be different at each occurrence. As usual, $e(z)=\exp(2\pi\mathrm{i}z)=\mathrm{e}^{2\pi\mathrm{i}z}$ and $\delta_{m,n}$ is the Kronecker symbol that detects $m=n$. We write $f=O(g)$ or $f\ll g$ to mean $|f|\leq cg$ for some unspecified positive constant $c$. The symbol $y\asymp Y$ is to mean that $c_{1}Y\leq|y|\leq c_{2}Y$ for some positive constants $c_{1}$ and $c_{2}$.
Structure of this paper. The rest of this paper is organized as follows. In Section 2, we give some preliminary results needed for the proof. In Section 3, we establish some lemmas and provide a crucial approximation of $S_{F}(t)$. In Section 4, we give the detailed proof of Theorem 1.1. In Section 5, we summarize our main results, highlight the significance of our work, and discuss prospects for future research.
2. Preliminaries
2.1. Hecke relations
Firstly, let us recall the following Hecke relations (see Goldfeld [10, pg. 168])
$\begin{split}A_{F}(m_{1}m_{2},n_{1}n_{2})&=A_{F}(m_{1},n_{1})A_{F}(m_{2},n_{2}),\quad\text{if}\,\,(m_{1}n_{1},m_{2}n_{2})=1,\\ A_{F}(n,1)A_{F}(m_{1},m_{2})&=\sum_{d_{0}d_{1}d_{2}=n\atop d_{1}|m_{1},\,d_{2}|m_{2}}A_{F}\bigg(\frac{m_{1}d_{0}}{d_{1}},\frac{m_{2}d_{1}}{d_{2}}\bigg),\\ A_{F}(1,n)A_{F}(m_{1},m_{2})&=\sum_{d_{0}d_{1}d_{2}=n\atop d_{1}|m_{1},\,d_{2}|m_{2}}A_{F}\bigg(\frac{m_{1}d_{2}}{d_{1}},\frac{m_{2}d_{0}}{d_{2}}\bigg),\\ A_{F}(n_{1},1)A_{F}(1,n_{2})&=\sum_{d|(n_{1},n_{2})}A_{F}\bigg(\frac{n_{1}}{d},\frac{n_{2}}{d}\bigg).\end{split}$
Then we arrive at the following useful result, which was presented by Liu and Liu in [15, Proposition 5.1].
Lemma 2.1.
For each prime $p$ and each integer $n\geq 0$, we have
$\begin{split}A_{F}(1,p)^{n}=\sum_{k,\ell\geq 0\atop k+\ell\leq n}a_{k,\ell}(n)A_{F}(p^{k},p^{\ell})\end{split}$
for some integers $a_{k,\ell}(n)\geq 0$. Furthermore, $a_{k,\ell}(n)\neq 0$ if and only if $(k,\ell)$ belongs to
$\begin{split}E_{n}:=\big\{(n_{2}-n_{1},n_{3}-n_{2})\,|\,n_{j}\in\mathbb{Z},n_{3}\geq n_{2}\geq n_{1}\geq 0,n_{1}+n_{2}+n_{3}=n\big\}.\end{split}$
2.2. A summation formula
We need the following summation formula, which is an application of the $\mathrm{GL}_{3}$ Kuznetsov trace formula in the version of Buttcane [6, Theorems 2–4].
Theorem 2.2.
Let $P=m_{1}n_{1}m_{2}n_{2}\neq 0$. For any $\varepsilon>0$, we have
$\begin{split}&\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\overline{A_{F}(m_{1},m_{2})}A_{F}(n_{1},n_{2})\\ &\quad=\delta_{n_{1},m_{1}}\delta_{n_{2},m_{2}}\mathcal{H}+O\bigg((TP)^{\varepsilon}\big(TP^{1/2}+T^{3}+TM^{2}P^{\vartheta}\big)\bigg),\end{split}$
where $\vartheta=7/64$, due to Kim and Sarnak [13], is the best record towards the Ramanujan–Petersson conjecture for Fourier coefficients of $\mathrm{GL}_{2}$ Maass cusp forms.
Proof.
See Sun and Wang [20, Theorem 7.1]. ∎
We quote here the frequently used estimate
$\mathcal{H}=CT^{3}M^{2}+O(T^{2}M^{3}),$
for some constant $C>0$ (see Buttcane and Zhou [7, pg. 15]). Taking $m_{1}=n_{1}=m_{2}=n_{2}=1$ in Theorem 2.2 immediately gives
Corollary 2.3.
Let the assumptions be as in Theorem 2.2. We have
$\begin{split}&\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}=CT^{3}M^{2}+O(T^{2}M^{3}).\end{split}$
2.3. The zero-density estimate
In the context of our proof, the key point lies in the following weighted version of zero-density estimate in the spectral aspect for the $L$-functions associated to $\mathrm{GL}_{3}$ Hecke–Maass cusp forms, which was rigorously established by Sun and Wang in [20, Theorem 1.1].
Lemma 2.4.
For $\sigma>0$ and $H>0$, we let
$\begin{split}N(\sigma,H;F):=\#\big\{\rho=1/2+\beta+\mathrm{i}\gamma:L(\rho,F)=0,\beta>\sigma,|\gamma|<H\big\}\end{split}$
and
$\begin{split}N(\sigma,H):=\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}N(\sigma,H;F).\end{split}$
Let $2/\log T<\sigma<1/2$. For some sufficiently small $\delta_{1},\,\theta>0$, we have
$\begin{split}N(\sigma,H)\ll HT^{-\theta\sigma}\log T,\end{split}$
uniformly in $3/\log T<H<T^{\delta_{1}}$.
3. An approximation of $S_{F}(t)$
In this section we prove several lemmas and derive an approximation of $S_{F}(t)$. We denote by $\rho=\beta+\mathrm{i}\gamma$ a typical (non-trivial) zero of $L(s,F)$ inside the critical strip, i.e., $0<\beta<1$. For $\operatorname{Re}(s)>1$, we have the Dirichlet series expansion for the logarithmic derivative of $L(s,F)$,
$\frac{L^{\prime}}{L}(s,F)=\sum_{n=1}^{\infty}\frac{\Lambda(n)C_{F}(n)}{n^{s}},$ (3.1)
where $\Lambda(n)$ denotes the von Mangoldt function, and
$C_{F}(n)=\begin{cases}\alpha_{F,1}(p)^{k}+\alpha_{F,2}(p)^{k}+\alpha_{F,3}(p)^{k},\,&\text{if}\,\,n=p^{k}\,\,\text{for a prime}\,\,p,\\ 0,\,&\text{otherwise}.\end{cases}$ (3.2)
Lemma 3.1.
Let $x>1$. For $s\neq\rho$ and $s\neq-2m+\mu_{F,j}$ ($m\geq 0$, $j=1,2,3$), we have the following identity
$\begin{split}\frac{L^{\prime}}{L}(s,F)=&-\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{s}}-\frac{1}{\log^{2}x}\sum_{\rho}\frac{x^{\rho-s}(1-x^{\rho-s})^{2}}{(\rho-s)^{3}}\\ &-\frac{1}{\log^{2}x}\sum_{j=1}^{3}\sum_{m=0}^{\infty}\frac{x^{-2m-s+\mu_{F,j}}(1-x^{-2m-s+\mu_{F,j}})^{2}}{(-2m-s+\mu_{F,j})^{3}},\end{split}$
where
[…]
Proof.
Recall the discontinuous integral
$\frac{1}{2\pi\mathrm{i}}\int_{(\alpha)}\frac{y^{s}}{s^{3}}\mathrm{d}s=\begin{cases}\frac{\log^{2}y}{2},\,&\textit{if }\,\,y\geq 1,\\ 0,\,&\textit{if }\,\,0<y\leq 1\end{cases}$
for $\alpha>0$. It follows from (3.1) and (3.2) that
$\log^{2}x\sum_{n=1}^{\infty}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{s}}=\frac{1}{2\pi\mathrm{i}}\int_{(\alpha)}\frac{x^{v}(1-x^{v})^{2}}{v^{3}}\frac{L^{\prime}}{L}(s+v,F)\mathrm{d}v,$
where $\alpha=\max\{2,1+\operatorname{Re}(s)\}$. By moving the line of integration all way to the left, we pick up the residues at $v=0$, $v=\rho-s$ and $v=-2m-s+\mu_{F,j}$ ($m\geq 0$, $j=1,2,3$), and deduce that
$\begin{split}&\frac{1}{2\pi\mathrm{i}}\int_{(\alpha)}\frac{x^{v}(1-x^{v})^{2}}{v^{3}}\frac{L^{\prime}}{L}(s+v,F)\mathrm{d}v\\ =&\frac{L^{\prime}}{L}(s,F)\log^{2}x+\sum_{\rho}\frac{x^{\rho-s}(1-x^{\rho-s})^{2}}{(\rho-s)^{3}}+\sum_{j=1}^{3}\sum_{m=0}^{\infty}\frac{x^{-2m-s+\mu_{F,j}}(1-x^{-2m-s+\mu_{F,j}})^{2}}{(-2m-s+\mu_{F,j})^{3}}.\end{split}$
This completes the proof of the lemma. ∎
Lemma 3.2.
For $s=\sigma+\mathrm{i}t$ and $s^{\prime}=\sigma^{\prime}+\mathrm{i}t^{\prime}$ such that $1/2\leq\sigma,\sigma^{\prime}\leq 10$, $s\neq\rho$ and $s^{\prime}\neq\rho$, we have
$\operatorname{Im}\bigg(\frac{L^{\prime}}{L}(s,F)-\frac{L^{\prime}}{L}(s^{\prime},F)\bigg)=\operatorname{Im}\sum_{\rho}\bigg(\frac{1}{s-\rho}-\frac{1}{s^{\prime}-\rho}\bigg)+O(1),$
and
$\operatorname{Re}\frac{L^{\prime}}{L}(s,F)=\sum_{\rho}\frac{\sigma-\beta}{(\sigma-\beta)^{2}+(t-\gamma)^{2}}+O(\log(|t|+M_{F})),$ (3.3)
where
$\begin{split}M_{F}:=2+\max_{1\leq j\leq 3}|\mu_{F,j}|.\end{split}$
Proof.
By the Hadamard factorization theorem for the entire function $\Lambda(s,F)$, we have
$\frac{L^{\prime}}{L}(s,F)=b_{F}+\sum_{\rho}\bigg(\frac{1}{s-\rho}+\frac{1}{\rho}\bigg)-\frac{1}{2}\sum_{j=1}^{3}\frac{\Gamma^{\prime}}{\Gamma}\bigg(\frac{s-\mu_{F,j}}{2}\bigg)+\frac{3}{2}\log\pi,$
for some $b_{F}\in\mathbb{C}$ with $\operatorname{Re}(b_{F})=-\operatorname{Re}\sum_{\rho}\frac{1}{\rho}$ (see [12, Proposition 5.7 (3)]). Note that for $z\not\in$-$\mathbb{N}$,
$\frac{\Gamma^{\prime}}{\Gamma}(z)=-\gamma+\sum_{k\geq 1}\bigg(\frac{1}{k}-\frac{1}{k-1+z}\bigg),$ (3.4)
where $\gamma$ is the Euler constant. Using (3.4), we get that for $\frac{1}{14}\leq u\leq\frac{145}{28}$ and $v\ll 1$,
$\operatorname{Im}\frac{\Gamma^{\prime}}{\Gamma}(u+\mathrm{i}v)=\sum_{k\geq 1}\frac{v}{(k+u-1)^{2}+v^{2}}\ll 1.$
Combining with the asymptotic formula of the logarithmic derivative of $\Gamma(z)$,
$\frac{\Gamma^{\prime}}{\Gamma}(z)=\log z-\frac{1}{2z}+O_{\varepsilon}\bigg(\frac{1}{|z|^{2}}\bigg),\quad\text{for}\,\,|\arg z|\leq\pi-\varepsilon,$
we complete the proof of the lemma. ∎
Let $x\geq 4$. We define
$\sigma_{x}=\sigma_{x,F}=\sigma_{x,F,t}:=\frac{1}{2}+2\max_{\rho}\bigg\{\bigg|\beta-\frac{1}{2}\bigg|,\frac{5}{\log x}\bigg\},$ (3.5)
where $\rho=\beta+\mathrm{i}\gamma$ runs through the zeros of $L(s,F)$ for which
$|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}.$
The following lemma is to display the dependence of $\sigma_{x}$ as needed.
Lemma 3.3.
Let $x\geq 4$. For $\sigma\geq\sigma_{x}$, we have
$\begin{split}\frac{L^{\prime}}{L}(\sigma+\mathrm{i}t,F)=&-\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma+\mathrm{i}t}}+O\bigg(x^{1/4-\sigma/2}\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|\bigg)\\ &+O\bigg(x^{1/4-\sigma/2}\log(|t|+M_{F})\bigg),\end{split}$
and
$\sum_{\rho}\frac{\sigma_{x}-1/2}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}=O\bigg(\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|\bigg)+O\big(\log(|t|+M_{F})\big).$
Proof.
By (3.3), we have
$\operatorname{Re}\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)=\sum_{\rho}\frac{\sigma_{x}-\beta}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}+O(\log(|t|+M_{F})).$ (3.6)
Moreover, if $\rho=\beta+\mathrm{i}\gamma$ is zero of $L(s,F)$, then $(1-\beta)+\mathrm{i}\gamma$ is also a zero of $L(s,F)$ by the functional equation (1.7). Thus we have
$\sum_{\rho}\frac{\sigma_{x}-\beta}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}$
$=$ $\,\frac{1}{2}\bigg(\sum_{\rho}\frac{\sigma_{x}-\beta}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}+\sum_{\rho}\frac{\sigma_{x}-(1-\beta)}{(\sigma_{x}-(1-\beta))^{2}+(t-\gamma)^{2}}\bigg)$
$=$ $\,\bigg(\sigma_{x}-\frac{1}{2}\bigg)\sum_{\rho}\frac{(\sigma_{x}-1/2)^{2}-(\beta-1/2)^{2}+(t-\gamma)^{2}}{((\sigma_{x}-\beta)^{2}+(t-\gamma)^{2})((\sigma_{x}-1+\beta)^{2}+(t-\gamma)^{2})}.$ (3.7)
If $|\beta-1/2|\leq\frac{\sigma_{x}-1/2}{2}$, then
$\begin{split}(\sigma_{x}-1/2)^{2}-(\beta-1/2)^{2}&\geq\frac{1}{2}\big((\sigma_{x}-1/2)^{2}+(\beta-1/2)^{2}\big)\\ &=\frac{1}{4}\big((\sigma_{x}-\beta)^{2}+(\sigma_{x}-1+\beta)^{2}\big).\end{split}$
Thus,
$(\sigma_{x}-1/2)^{2}-(\beta-1/2)^{2}+(t-\gamma)^{2}\geq\frac{1}{4}\big((\sigma_{x}-1+\beta)^{2}+(t-\gamma)^{2}\big).$ (3.8)
If $|\beta-1/2|>\frac{\sigma_{x}-1/2}{2}$, then by (3.5) we have
$|t-\gamma|>\frac{x^{3|\beta-1/2|}}{\log x}>3|\beta-1/2|.$
Thus,
$(\sigma_{x}-1/2)^{2}-(\beta-1/2)^{2}+(t-\gamma)^{2}$
$=\big((\sigma_{x}-1/2)^{2}+(\beta-1/2)^{2}\big)+(t-\gamma)^{2}-2(\beta-1/2)^{2}$
$\geq\frac{1}{2}\big((\sigma_{x}-\beta)^{2}+(\sigma_{x}-1+\beta)^{2}\big)+\frac{7}{9}(t-\gamma)^{2}$
$\geq\frac{1}{4}\big((\sigma_{x}-1+\beta)^{2}+(t-\gamma)^{2}\big).$ (3.9)
Combining (3.8), (3) and (3), we get
$\sum_{\rho}\frac{\sigma_{x}-\beta}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}\geq\frac{1}{4}(\sigma_{x}-1/2)\sum_{\rho}\frac{1}{((\sigma_{x}-\beta)^{2}+(t-\gamma)^{2})}.$
Using this bound and (3.6), we obtain
$\sum_{\rho}\frac{1}{((\sigma_{x}-\beta)^{2}+(t-\gamma)^{2})}\leq\frac{4}{\sigma_{x}-1/2}\bigg|\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)\bigg|+O\bigg(\frac{\log(|t|+M_{F})}{\sigma_{x}-1/2}\bigg).$ (3.10)
On the other hand, by Lemma 3.1 and the estimate (1.6), we have, for $\sigma\geq\sigma_{x}$,
$\frac{L^{\prime}}{L}(\sigma+\mathrm{i}t,F)=$ $-\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma+\mathrm{i}t}}+\frac{w_{F}(x,\sigma,t)}{\log^{2}x}\sum_{\rho}\frac{x^{\beta-\sigma}(1+x^{\beta-\sigma})^{2}}{\big((\sigma-\beta)^{2}+(t-\gamma)^{2}\big)^{3/2}}$
$+O\bigg(\frac{x^{5/14-\sigma}}{\log^{2}x}\bigg),$ (3.11)
with $|w_{F}(x,\sigma,t)|\leq 1$. Next we claim that
$\frac{x^{\beta-\sigma}(1+x^{\beta-\sigma})^{2}}{\big((\sigma-\beta)^{2}+(t-\gamma)^{2}\big)^{3/2}}\leq 2\log x\frac{x^{1/4-\sigma/2}}{(\beta-\sigma_{x})^{2}+(t-\gamma)^{2}}.$ (3.12)
If $\beta\leq\frac{\sigma_{x}+1/2}{2}$, then we have
$\begin{split}\frac{x^{\beta-\sigma}(1+x^{\beta-\sigma})^{2}}{\big((\sigma-\beta)^{2}+(t-\gamma)^{2}\big)^{3/2}}&\leq\frac{4x^{1/4-\sigma/2}}{(\sigma_{x}-\beta)\big((\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}\big)}\\ &\leq\frac{8}{\sigma_{x}-1/2}\frac{x^{1/4-\sigma/2}}{\big((\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}\big)}\\ &\leq\frac{4}{5}\log x\frac{x^{1/4-\sigma/2}}{\big((\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}\big)}.\end{split}$
If $\beta>\frac{\sigma_{x}+1/2}{2}$, then we have
$|t-\gamma|>\frac{x^{3|\beta-1/2|}}{\log x}>3|\beta-1/2|\geq 3|\beta-\sigma_{x}|.$
Thus, $(t-\gamma)^{2}>\frac{8}{9}\big((\beta-\sigma_{x})^{2}+(t-\gamma)^{2}\big)$. Hence
$\begin{split}\frac{x^{\beta-\sigma}(1+x^{\beta-\sigma})^{2}}{\big((\sigma-\beta)^{2}+(t-\gamma)^{2}\big)^{3/2}}&\leq\frac{x^{\beta-\sigma}(1+x^{\beta-1/2})^{2}}{|t-\gamma|(t-\gamma)^{2}}\leq\frac{\log x}{x^{3|\beta-1/2|}}\frac{9}{8}\frac{x^{\beta-\sigma}(1+x^{\beta-1/2})^{2}}{(\beta-\sigma_{x})^{2}+(t-\gamma)^{2}}\\ &=\frac{9}{8}(\log x)(1+x^{-(\beta-1/2)})^{2}\frac{x^{1/2-\sigma}}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}\\ &\leq\frac{9}{8}(\log x)(1+e^{-5})^{2}\frac{x^{1/2-\sigma}}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}\\ &<2(\log x)\frac{x^{1/2-\sigma}}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}.\end{split}$
Using (3.10) and (3.12), we have
$\begin{split}&\sum_{\rho}\frac{x^{\beta-\sigma}(1+x^{\beta-\sigma})^{2}}{\big((\sigma-\beta)^{2}+(t-\gamma)^{2}\big)^{3/2}}\\ \leq&\,2\log x\sum_{\rho}\frac{x^{1/4-\sigma/2}}{(\beta-\sigma_{x})^{2}+(t-\gamma)^{2}}\\ \leq&\,8\log x\frac{x^{1/4-\sigma/2}}{\sigma_{x}-1/2}\bigg|\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)\bigg|+O\bigg(\frac{(\log x)x^{1/4-\sigma/2}\log(|t|+M_{F})}{\sigma_{x}-1/2}\bigg)\\ \leq&\,\frac{4}{5}(\log x)^{2}x^{1/4-\sigma/2}\bigg|\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)\bigg|+O\bigg((\log x)^{2}x^{1/4-\sigma/2}\log\big(|t|+M_{F}\big)\bigg).\end{split}$
Inserting this bound into (3), we have
$\frac{L^{\prime}}{L}(\sigma+\mathrm{i}t,F)=$ $-\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma+\mathrm{i}t}}+\frac{4}{5}w^{\prime}_{F}(x,\sigma,t)x^{1/4-\sigma/2}\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)$
$+O\bigg(x^{1/4-\sigma/2}\log(|t|+M_{F})\bigg),$ (3.13)
with $|w^{\prime}_{F}(x,\sigma,t)|\leq 1$. By taking $\sigma=\sigma_{x}$,
$\begin{split}&\bigg(1-\frac{4}{5}w^{\prime}_{F}(x,\sigma_{x},t)x^{1/4-\sigma_{x}/2}\bigg)\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)\\ =&\,O\bigg(\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|\bigg)+O\bigg(x^{1/4-\sigma_{x}/2}\log(|t|+M_{F})\bigg).\end{split}$
Since $\Big|1-\frac{4}{5}w^{\prime}_{F}(x,\sigma_{x},t)x^{1/4-\sigma_{x}/2}\Big|\geq 1-\Big|\frac{4}{5}w^{\prime}_{F}(x,\sigma_{x},t)x^{1/4-\sigma_{x}/2}\Big|\geq\frac{1}{5}$, we have
$\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)=O\bigg(\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|\bigg)+O\bigg(x^{1/4-\sigma_{x}/2}\log(|t|+M_{F})\bigg).$ (3.14)
Then the claimed results follow from (3) and (3.14), and from (3.10) and (3.14), respectively. ∎
The following theorem is to provide an approximation of $S_{F}(t)$.
Theorem 3.4.
For $t\neq 0$ and $x\geq 4$, we have
$\begin{split}S_{F}(t)=&\frac{1}{\pi}\operatorname{Im}\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}\log n}+O\bigg((\sigma_{x}-1/2)\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|\bigg)\\ &+O\big((\sigma_{x}-1/2)\log(|t|+M_{F})\big),\end{split}$
where $\sigma_{x}$ is defined in (3.5).
Proof.
By the definition of $S_{F}(t)$, we have
$\begin{split}\pi S_{F}(t)=&-\int_{1/2}^{\infty}\operatorname{Im}\frac{L^{\prime}}{L}(\sigma+\mathrm{i}t,F)\mathrm{d}\sigma\\ =&-\int_{\sigma_{x}}^{\infty}\operatorname{Im}\frac{L^{\prime}}{L}(\sigma+\mathrm{i}t,F)\mathrm{d}\sigma-(\sigma_{x}-1/2)\operatorname{Im}\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)\\ &+\int_{1/2}^{\sigma_{x}}\operatorname{Im}\bigg(\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)-\frac{L^{\prime}}{L}(\sigma+\mathrm{i}t,F)\bigg)\mathrm{d}\sigma\\ :=&\,J_{1}+J_{2}+J_{3}.\end{split}$
By Lemma 3.3, we have
$\begin{split}J_{1}=&-\int_{\sigma_{x}}^{\infty}\operatorname{Im}\frac{L^{\prime}}{L}(\sigma+\mathrm{i}t,F)\mathrm{d}\sigma\\ =&\int_{\sigma_{x}}^{\infty}\operatorname{Im}\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma+\mathrm{i}t}}\mathrm{d}\sigma+O\bigg(\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|\int_{\sigma_{x}}^{\infty}x^{1/4-\sigma/2}\mathrm{d}\sigma\bigg)\\ &+O\bigg(\log(|t|+M_{F})\int_{\sigma_{x}}^{\infty}x^{1/4-\sigma/2}\mathrm{d}\sigma\bigg)\\ =&\,\operatorname{Im}\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}\log n}+O\bigg(\frac{1}{\log x}\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|\bigg)+O\bigg(\frac{\log(|t|+M_{F})}{\log x}\bigg).\end{split}$
Taking $\sigma=\sigma_{x}$ in Lemma 3.3, we have
$\begin{split}J_{2}\leq&\,(\sigma_{x}-1/2)\bigg|\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)\bigg|\\ \ll&\,(\sigma_{x}-1/2)\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|+(\sigma_{x}-1/2)\log(|t|+M_{F}).\end{split}$
Using Lemma 3.2, we get
$\begin{split}&\operatorname{Im}\bigg(\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)-\frac{L^{\prime}}{L}(\sigma+\mathrm{i}t,F)\bigg)\\ =&\sum_{\rho}\frac{(\sigma_{x}-\sigma)(\sigma+\sigma_{x}-2\beta)(t-\gamma)}{\big((\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}\big)\big((\sigma-\beta)^{2}+(t-\gamma)^{2}\big)}+O(1).\end{split}$
Hence
$\begin{split}|J_{3}|\leq&\sum_{\rho}\int_{1/2}^{\sigma_{x}}\frac{(\sigma_{x}-\sigma)|\sigma+\sigma_{x}-2\beta||t-\gamma|}{\big((\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}\big)\big((\sigma-\beta)^{2}+(t-\gamma)^{2}\big)}\mathrm{d}\sigma+O(\sigma_{x}-1/2)\\ \leq&\sum_{\rho}\frac{\sigma_{x}-1/2}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}\int_{1/2}^{\sigma_{x}}\frac{|\sigma+\sigma_{x}-2\beta||t-\gamma|}{(\sigma-\beta)^{2}+(t-\gamma)^{2}}\mathrm{d}\sigma+O(\sigma_{x}-1/2).\end{split}$
If $|\beta-1/2|\leq\frac{1}{2}(\sigma_{x}-1/2)$, then for $1/2\leq\sigma\leq\sigma_{x}$,
$\begin{split}|\sigma+\sigma_{x}-2\beta|&=|(\sigma-1/2)+(\sigma_{x}-1/2)-2(\beta-1/2)|\\ &\leq|\sigma-1/2|+|\sigma_{x}-1/2|+2|\beta-1/2|\leq 3(\sigma_{x}-1/2).\end{split}$
Thus,
$\int_{1/2}^{\sigma_{x}}\frac{|\sigma+\sigma_{x}-2\beta||t-\gamma|}{(\sigma-\beta)^{2}+(t-\gamma)^{2}}\mathrm{d}\sigma$ $\leq 3(\sigma_{x}-1/2)\int_{-\infty}^{\infty}\frac{|t-\gamma|}{(\sigma-\beta)^{2}+(t-\gamma)^{2}}\mathrm{d}\sigma$
$\leq 10(\sigma_{x}-1/2).$ (3.15)
If $|\beta-1/2|>\frac{1}{2}(\sigma_{x}-1/2)$, then
$|t-\gamma|>\frac{x^{3|\beta-1/2|}}{\log x}>3|\beta-1/2|$
and for $1/2\leq\sigma\leq\sigma_{x}$,
$\begin{split}|\sigma+\sigma_{x}-2\beta|\leq|\sigma-1/2|+|\sigma_{x}-1/2|+2|\beta-1/2|\leq 6|\beta-1/2|.\end{split}$
Thus,
$\int_{1/2}^{\sigma_{x}}\frac{|\sigma+\sigma_{x}-2\beta||t-\gamma|}{(\sigma-\beta)^{2}+(t-\gamma)^{2}}\mathrm{d}\sigma$ $<\int_{1/2}^{\sigma_{x}}\frac{|\sigma+\sigma_{x}-2\beta|}{|t-\gamma|}\mathrm{d}\sigma$
$\leq\int_{1/2}^{\sigma_{x}}\frac{6|\beta-1/2|}{3|\beta-1/2|}\mathrm{d}\sigma=2(\sigma_{x}-1/2).$ (3.16)
By the estimates in (3)–(3) and Lemma 3.3, we obtain
$\begin{split}|J_{3}|\leq&10(\sigma_{x}-1/2)\sum_{\rho}\frac{\sigma_{x}-1/2}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}+O(\sigma_{x}-1/2)\\ =&O\bigg((\sigma_{x}-1/2)\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|\bigg)+O\big((\sigma_{x}-1/2)\log(|t|+M_{F})\big).\end{split}$
Thus we complete the proof of the theorem. ∎
4. Proof of Theorem 1.1
Lemma 4.1.
Let $t>0$ be fixed and $T$ be any large positive parameter. For $k\in\mathbb{N}$, $x=T^{\delta/3}$ with $0<\delta<\frac{3\theta}{8nk+3}$, we have
$\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{4n}x^{4nk(\sigma_{x}-1/2)}\ll_{t,n,\delta,k}\frac{1}{(\log T)^{4n}},\end{split}$
where $\sigma_{x,F}$ is defined in (3.5) and $\theta$ is as in Lemma 2.4.
Proof.
For $k\in\mathbb{N}$, by the definition of $\sigma_{x,F}$, we have
$\begin{split}&(\sigma_{x,F}-1/2)^{4n}x^{4nk(\sigma_{x,F}-1/2)}\\ \leq&\bigg(\frac{10}{\log x}\bigg)^{4n}x^{40nk/\log x}+2^{4n+1}\sum_{\beta>\frac{1}{2}+\frac{5}{\log x}\atop|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}}(\beta-1/2)^{4n}x^{8nk(\beta-1/2)}.\end{split}$
On the other hand,
$\begin{split}&\sum_{\beta>\frac{1}{2}+\frac{5}{\log x}\atop|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}}(\beta-1/2)^{4n}x^{8nk(\beta-1/2)}\\ \leq&\sum_{i=5}^{\frac{1}{2}\lfloor\log x\rfloor}\bigg(\frac{i+1}{\log x}\bigg)^{4n}x^{8nk\frac{i+1}{\log x}}\sum_{\frac{1}{2}+\frac{i}{\log x}<\beta\leq\frac{1}{2}+\frac{i+1}{\log x}\atop|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}}1\\ \leq&\frac{1}{(\log x)^{4n}}\sum_{i=5}^{\frac{1}{2}\lfloor\log x\rfloor}\big(i+1\big)^{4n}\mathrm{e}^{8nk(i+1)}N\bigg(\frac{i}{\log x},|t|+\frac{\mathrm{e}^{3(i+1)}}{\log x};F\bigg).\end{split}$
By Lemma 2.4, we have
$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\sum_{\beta>\frac{1}{2}+\frac{5}{\log x}\atop|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}}(\beta-1/2)^{4n}x^{8nk(\beta-1/2)}\\ \leq&\,\frac{1}{(\log x)^{4n}}\sum_{i=5}^{\frac{1}{2}\lfloor\log x\rfloor}\big(i+1\big)^{4n}\mathrm{e}^{8nk(i+1)}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}N\bigg(\frac{i}{\log x},|t|+\frac{\mathrm{e}^{3(i+1)}}{\log x};F\bigg)\\ \ll&\,\frac{1}{(\log x)^{4n}}\sum_{i=5}^{\infty}\big(i+1\big)^{4n}\mathrm{e}^{8nk(i+1)}\bigg(|t|+\frac{\mathrm{e}^{3(i+1)}}{\log x}\bigg)T^{-\theta\frac{i}{\log x}}\log T\\ \ll&{}_{t,n,\delta,k}\,\frac{\log T}{(\log x)^{4n+1}}\sum_{i=5}^{\infty}\big(i+1\big)^{4n}\mathrm{e}^{(8nk+3-\frac{3\theta}{\delta})i}\\ \ll&{}_{t,n,\delta,k}\,\frac{1}{(\log T)^{4n}}\end{split}$
provided that $0<\delta<\frac{3\theta}{8nk+3}$. In addition, by Corollary 2.3 and (1.9),
$\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg(\frac{10}{\log x}\bigg)^{4n}x^{40nk/\log x}\ll_{n,k}\frac{1}{(\log x)^{4n}}\ll_{n,k,\delta}\frac{1}{(\log T)^{4n}}.\end{split}$
Thus we complete the proof of the lemma. ∎
For a positive parameter $x$ (to be chosen later), let
$\mathcal{M}_{F}(t):=\frac{1}{\pi}\operatorname{Im}\sum_{p\leq x^{3}}\frac{C_{F}(p)}{p^{1/2+\mathrm{i}t}}\qquad\text{and}\qquad\mathcal{R}_{F}(t):=S_{F}(t)-\mathcal{M}_{F}(t),$ (4.1)
where we recall (3.2) that $C_{F}(p)=\alpha_{F,1}(p)+\alpha_{F,2}(p)+\alpha_{F,3}(p)=A_{F}(1,p)$.
Lemma 4.2.
We have
$\begin{split}\mathcal{R}_{F}(t)=&O\bigg(\bigg|\operatorname{Im}\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+\mathrm{i}t}\log p}\bigg|\bigg)+O\bigg(\sum_{k=2}^{7}\bigg|\operatorname{Im}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+\mathrm{i}t)}\log p}\bigg|\bigg)\\ &+O\bigg((\sigma_{x}-1/2)\sum_{k=1}^{7}x^{k(\sigma_{x}-1/2)}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma\bigg)\\ &+O\big((\sigma_{x}-1/2)\log(|t|+M_{F})\big)+O\big(1\big).\end{split}$
Proof.
By Theorem 3.4,
$\begin{split}\mathcal{R}_{F}(t)=&\frac{1}{\pi}\operatorname{Im}\sum_{p\leq x^{3}}\frac{C_{F}(p)(\Lambda_{x}(p)p^{1/2-\sigma_{x}}-\Lambda(p))}{p^{1/2+\mathrm{i}t}\log p}+\frac{1}{\pi}\operatorname{Im}\sum_{k=2}^{\infty}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+\mathrm{i}t)}\log p^{k}}\\ &+O\bigg((\sigma_{x}-1/2)\bigg|\sum_{k=1}^{\infty}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+\mathrm{i}t)}}\bigg|\bigg)+O\big((\sigma_{x}-1/2)\log(|t|+M_{F})\big).\end{split}$
Using the bound (1.6), we deduce that
$\begin{split}\sum_{k=8}^{\infty}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+\mathrm{i}t)}}=O(1),\quad\text{and}\quad\sum_{k=8}^{\infty}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+\mathrm{i}t)}\log p}=O(1).\end{split}$
Thus,
$\begin{split}\mathcal{R}_{F}(t)=&\,\frac{1}{\pi}\operatorname{Im}\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+\mathrm{i}t}\log p}+\frac{1}{\pi}\operatorname{Im}\sum_{k=2}^{7}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+\mathrm{i}t)}\log p^{k}}\\ &+\frac{1}{\pi}\operatorname{Im}\sum_{k=1}^{7}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})(p^{k(1/2-\sigma_{x})}-1)}{p^{k(1/2+\mathrm{i}t)}\log p^{k}}\\ &+O\bigg((\sigma_{x}-1/2)\bigg|\sum_{k=1}^{7}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+\mathrm{i}t)}}\bigg|\bigg)\\ &+O\big((\sigma_{x}-1/2)\log(|t|+M_{F})\big)+O(1).\end{split}$
Note that for $1/2\leq a\leq\sigma_{x}$ and $k\geq 1$,
$\begin{split}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(a+\mathrm{i}t)}}\bigg|&=kx^{k(a-1/2)}\bigg|\int_{a}^{\infty}x^{k(1/2-\sigma)}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\mathrm{d}\sigma\bigg|\\ &\leq kx^{k(\sigma_{x}-1/2)}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma.\end{split}$
Thus,
$\begin{split}&\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})(p^{k(1/2-\sigma_{x})}-1)}{p^{k(1/2+\mathrm{i}t)}\log p^{k}}\bigg|=\bigg|\int_{1/2}^{\sigma_{x}}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(a+\mathrm{i}t)}}\mathrm{d}a\bigg|\\ &\leq k(\sigma_{x}-1/2)x^{k(\sigma_{x}-1/2)}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma,\end{split}$
and
$\begin{split}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+\mathrm{i}t)}}\bigg|\leq kx^{k(\sigma_{x}-1/2)}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma.\end{split}$
Thus we complete the proof of the lemma. ∎
Lemma 4.3.
Let $t>0$ be fixed and $T$ be any large positive parameter. For $x=T^{\delta/3}$ with sufficiently small $\delta>0$, we have
$\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\mathcal{M}_{F}(t)^{n}=C_{n}(\log\log T)^{\frac{n}{2}}+O_{t,n}\big((\log\log T)^{\frac{n-1}{2}}\big)\end{split}$ (4.2)
and
$\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{R}_{F}(t)|^{2n}=O_{t,n}(1),\end{split}$ (4.3)
where the $O$-term depends only on $t$ and $n$, $C_{n}$ is defined by
$C_{n}=\begin{cases}\frac{n!}{(n/2)!(2\pi)^{n}},\,&\textit{if }\,n\,\,\text{is even},\\ 0,\,&\textit{if }\,n\,\,\text{is odd}.\end{cases}$
Proof.
Recall that
$\mathcal{M}_{F}(t)=\frac{1}{\pi}\operatorname{Im}\sum_{p\leq x^{3}}\frac{C_{F}(p)}{p^{1/2+\mathrm{i}t}}=\frac{-\mathrm{i}}{2\pi}\bigg(\sum_{p\leq x^{3}}\frac{A_{F}(1,p)}{p^{1/2+\mathrm{i}t}}-\sum_{p\leq x^{3}}\frac{\overline{A_{F}(1,p)}}{p^{1/2-\mathrm{i}t}}\bigg).$
Set $x=T^{\delta/3}$ for a suitably small $\delta>0$. Thus we have that the general term of
$\mathcal{M}_{F}(t)^{n}=\frac{(-\mathrm{i})^{n}}{(2\pi)^{n}}\bigg(\sum_{p\leq T^{\delta}}\frac{A_{F}(1,p)}{p^{1/2+\mathrm{i}t}}-\sum_{p\leq T^{\delta}}\frac{\overline{A_{F}(1,p)}}{p^{1/2-\mathrm{i}t}}\bigg)^{n}$
has the form
$\frac{A_{F}(1,p_{1})^{m(p_{1})}\overline{A_{F}(1,p_{1})^{n(p_{1})}}}{p_{1}^{m(p_{1})(1/2+\mathrm{i}t)}(-1)^{n(p_{1})}p_{1}^{n(p_{1})(1/2-\mathrm{i}t)}}\times\cdots\times\frac{A_{F}(1,p_{r})^{m(p_{r})}\overline{A_{F}(1,p_{r})^{n(p_{r})}}}{p_{r}^{m(p_{r})(1/2+\mathrm{i}t)}(-1)^{n(p_{r})}p_{r}^{n(p_{r})(1/2-\mathrm{i}t)}},$ (4.4)
where $p_{1}<p_{2}<\cdots<p_{r}\leq T^{\delta}$, $m(p_{j})+n(p_{j})\geq 1$, and $\sum_{j=1}^{r}(m(p_{j})+n(p_{j}))=n$.
Now we discuss the contribution from the general term (4.4) in the following three cases.
Case (I): In the general term (4.4), $m(p_{j_{0}})\not\equiv n(p_{j_{0}})(\bmod\,3)$ for some $j_{0}$. By Lemma 2.1, we have
$\begin{split}A_{F}(1,p_{j_{0}})^{m(p_{j_{0}})}=\sum_{k,\ell\geq 0\atop k+\ell\leq m(p_{j_{0}})}a_{k,\ell}(m(p_{j_{0}}))A_{F}(p_{j_{0}}^{k},p_{j_{0}}^{\ell})\end{split}$
and
$\begin{split}\overline{A_{F}(1,p_{j_{0}})^{n(p_{j_{0}})}}=\sum_{k,\ell\geq 0\atop k+\ell\leq n(p_{j_{0}})}a_{k,\ell}(n(p_{j_{0}}))\overline{A_{F}(p_{j_{0}}^{k},p_{j_{0}}^{\ell})}.\end{split}$
Then we can observe that there does not exist $(k,\ell)$ such that $a_{k,\ell}(m(p_{j_{0}}))\neq 0$ and $a_{k,\ell}(n(p_{j_{0}}))\neq 0$. Otherwise, by Lemma 2.1, there would exist $(k,\ell)\in E_{m(p_{j_{0}})}\cap E_{n(p_{j_{0}})}$ and integers $m_{h},\,n_{h}\geq 0$, $h=1,2,3$, such that
$\begin{split}m(p_{j_{0}})&=m_{1}+m_{2}+m_{3},\\ n(p_{j_{0}})&=n_{1}+n_{2}+n_{3},\\ k&=-m_{1}+m_{2}=-n_{1}+n_{2},\\ \ell&=-m_{2}+m_{3}=-n_{2}+n_{3}.\end{split}$
There equations would further lead to
$\begin{split}m(p_{j_{0}})-n(p_{j_{0}})&=(m_{1}-n_{1})+(m_{2}-n_{2})+(m_{3}-n_{3})\\ &=3(m_{2}-n_{2})\equiv 0(\bmod\,3),\end{split}$
which contradicts the assumption for Case (I).
With the above observation and the Hecke relations, the numerator of this general term (4.4) can be written as a sum whose terms are all of the form $A_{F}(m_{1},n_{1})\overline{A_{F}(m_{2},n_{2})}$ where $(m_{1},n_{1})\neq(m_{2},n_{2})$. Since for these terms $\delta_{m_{1},m_{2}}\delta_{n_{1},n_{2}}=0$, we see from Theorem 2.2 and the estimate (1.9) that the contribution from the general term (4.4) to the moment (4.2) is
$\begin{split}&\ll_{n}\frac{1}{\mathcal{H}}\sum_{p_{1}<p_{2}<\cdots<p_{r}\leq T^{\delta}}p_{1}^{-\frac{1}{2}(m(p_{1})+n(p_{1}))}\cdots p_{r}^{-\frac{1}{2}(m(p_{r})+n(p_{r}))}\\ &\qquad\qquad\qquad\qquad\qquad\times\bigg((T^{1+2n\delta})^{\varepsilon}\big(T^{1+n\delta}+T^{3}+T^{1+2n\delta\vartheta}M^{2}\big)\bigg)\\ &\ll_{n}\frac{1}{\mathcal{H}}\bigg(\sum_{p\leq T^{\delta}}p^{-\frac{1}{2}}\bigg)^{r}\big(T^{1+\varepsilon+n\delta(1+2\varepsilon)}+T^{3+\varepsilon+2n\delta\varepsilon}+T^{1+\varepsilon+2n\delta(\vartheta+\varepsilon)}M^{2}\big)\\ &\ll_{n}T^{-2+\varepsilon+n\delta(3/2+2\varepsilon)}M^{-2}+T^{\varepsilon+n\delta(1/2+2\varepsilon)}M^{-2}+T^{-2+\varepsilon+n\delta(1/2+2\vartheta+2\varepsilon)},\end{split}$
which is negligible upon taking $0<\delta<\frac{1}{n}$.
Case (II): In the general term (4.4), $m(p_{j})\equiv n(p_{j})(\bmod\,3)$ for all $j$, and $m(p_{j_{0}})\geq 2$ or $n(p_{j_{0}})\geq 2$ for some $j_{0}$. In this case, $m(p_{j})+n(p_{j})\geq 2$ for all $j$ and $m(p_{j_{0}})+n(p_{j_{0}})\geq 3$ for some $j_{0}$, so $n>2r$ and $r<\lfloor\frac{n-1}{2}\rfloor$. By applying Lemma 2.1 to the numerator of this general term (4.4) and multiplying out by Hecke relations, one sees that it is possible to have resulting terms $A_{F}(m_{1},n_{1})\overline{A_{F}(m_{2},n_{2})}$ where $(m_{1},n_{1})=(m_{2},n_{2})$. Using Theorem 2.2, the contribution to the moment (4.2) such terms is bounded by the diagonal term of size $\asymp T^{3}M^{2}$. The contribution from the off-diagonal terms is clearly negligible compared to $T^{3}M^{2}$, given that $0<\delta<\frac{1}{n}$. Using the classical result
$\sum_{p\leq x}\frac{1}{p}=\log\log x+O(1),$ (4.5)
we have the contribution from this general term to the moment (4.2) is at most
$\begin{split}&\ll_{n}\sum_{p_{1}<p_{2}<\cdots<p_{r}\leq T^{\delta}}p_{1}^{-\frac{1}{2}(m(p_{1})+n(p_{1}))}\cdots p_{r}^{-\frac{1}{2}(m(p_{r})+n(p_{r}))}\\ &\ll_{n}\bigg(\sum_{p\leq T^{\delta}}p^{-1}\bigg)^{r}\ll_{n}(\log\log T)^{\lfloor\frac{n-1}{2}\rfloor}.\end{split}$
It remains to discuss the following: In the general term (4.4), $m(p_{j})=n(p_{j})=1$ for all $j$. Clearly $n$ must be even, say $n=2m$ and so $r=m$.
Case (III): $m(p_{j})=n(p_{j})=1$ for all $j$. The contribution from these terms to (4.2) is
$\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\binom{2m}{m}(m!)(m!)\frac{(-1)^{m}(-\mathrm{i})^{2m}}{(2\pi)^{2m}}$
$\qquad\sum_{p_{1}<p_{2}<\cdots<p_{m}\leq T^{\delta}}\frac{A_{F}(1,p_{1})\overline{A_{F}(1,p_{1})}\cdots A_{F}(1,p_{m})\overline{A_{F}(1,p_{m})}}{p_{1}\cdots p_{m}}.$ (4.6)
By the Hecke relations, we have
$\begin{split}&A_{F}(1,p_{1})\overline{A_{F}(1,p_{1})}\cdots A_{F}(1,p_{m})\overline{A_{F}(1,p_{m})}\\ &\quad=A_{F}(1,p_{1}\cdots p_{m})\overline{A_{F}(1,p_{1}\cdots p_{m})}.\end{split}$
Applying Theorem 2.2 and the estimate (4.5), we have (4) equals
$\begin{split}&\frac{1}{\mathcal{H}}\frac{(2m)!}{(2\pi)^{2m}}\sum_{p_{1}<p_{2}<\cdots<p_{m}\leq T^{\delta}}\frac{1}{p_{1}\cdots p_{m}}\cdot\bigg(\mathcal{H}+O\bigg(T^{1+\varepsilon}(p_{1}\cdots p_{m})+T^{3+\varepsilon}+T^{1+\varepsilon}M^{2}(p_{1}\cdots p_{m})^{2\vartheta}\bigg)\bigg)\\ =&\frac{(2m)!}{m!(2\pi)^{2m}}\sum_{p_{1},p_{2},\ldots,p_{m}\leq T^{\delta}\atop p_{j}\,\text{distinct}}\frac{1}{p_{1}\cdots p_{m}}+O\bigg(T^{-2+\varepsilon}M^{-2}\sum_{p_{1},p_{2},\ldots,p_{m}\leq T^{\delta}\atop p_{j}\,\text{distinct}}1\bigg)\\ &+O\bigg(T^{\varepsilon}M^{-2}\sum_{p_{1},p_{2},\ldots,p_{m}\leq T^{\delta}\atop p_{j}\,\text{distinct}}\frac{1}{p_{1}\cdots p_{m}}\bigg)+O\bigg(T^{-2+\varepsilon}\sum_{p_{1},p_{2},\ldots,p_{m}\leq T^{\delta}\atop p_{j}\,\text{distinct}}\frac{1}{(p_{1}\cdots p_{m})^{1-2\vartheta}}\bigg)\\ =&\frac{(2m)!}{m!(2\pi)^{2m}}(\log\log T+O(1))^{m}+O\big(T^{m\delta-2+\varepsilon}M^{-2}+T^{\varepsilon}M^{-2}(\log\log T)^{m}+T^{2m\delta\vartheta-2+\varepsilon}(\log\log T)^{m}\big)\\ =&\frac{(2m)!}{m!(2\pi)^{2m}}(\log\log T)^{m}+O_{m}\big((\log\log T)^{m-1}\big).\end{split}$
Now the asymptotic formula (4.2) follows from Cases (I)–(III).
Next we are ready to prove (4.3). Recall that $\sigma_{x}=\sigma_{x,F}$ depending on $F$. By Lemma 4.2,
$\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{R}_{F}(t)|^{2n}$
$\ll$ $\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg|\operatorname{Im}\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+\mathrm{i}t}\log p}\bigg|^{2n}+\sum_{k=2}^{7}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg|\operatorname{Im}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+\mathrm{i}t)}\log p}\bigg|^{2n}$
$+$ $\sum_{k=1}^{7}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{2n}x^{2nk(\sigma_{x}-1/2)}\bigg(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma\bigg)^{2n}$
$+$ $\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{2n}\big(\log(|t|+M_{F})\big)^{2n}+\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}.$ (4.7)
For the first two terms in (4), we have
$\begin{split}&\bigg|\operatorname{Im}\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+\mathrm{i}t}\log p}\bigg|^{2n}\\ &\qquad\ll_{n}\bigg(\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+\mathrm{i}t}\log p}-\sum_{p\leq x^{3}}\frac{\overline{A_{F}(1,p)}(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2-\mathrm{i}t}\log p}\bigg)^{2n},\end{split}$
and for $k\geq 2$,
$\begin{split}\bigg|\operatorname{Im}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+\mathrm{i}t)}\log p}\bigg|^{2n}\ll_{n}\bigg(\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+\mathrm{i}t)}\log p}-\sum_{p^{k}\leq x^{3}}\frac{\overline{C_{F}(p^{k})}\Lambda_{x}(p^{k})}{p^{k(1/2-\mathrm{i}t)}\log p}\bigg)^{2n}.\end{split}$ (4.8)
For (4.8), we note that
$\begin{split}C_{F}(p^{k})=\alpha_{F,1}(p)^{k}+\alpha_{F,2}(p)^{k}+\alpha_{F,3}(p)^{k}\end{split}$
is a symmetric polynomial in $\alpha_{F,j}(p)$, since (1.5) says that $1$, $A_{F}(1,p)$, $A_{F}(p,1)=\overline{A_{F}(1,p)}$ are the elementary symmetric polynomials in $\alpha_{F,j}(p)$, the fundamental theorem of symmetric polynomials (see [17, pg. 21, Remark]) implies that the symmetric polynomial $C_{F}(p^{k})=\alpha_{F,1}(p)^{k}+\alpha_{F,2}(p)^{k}+\alpha_{F,3}(p)^{k}$ is a polynomial in $1$, $A_{F}(1,p)$ and $\overline{A_{F}(1,p)}$. Then we can use the fact $|\Lambda_{x}(p)-\Lambda(p)|=O\big(\frac{\log^{3}p}{\log^{2}x}\big)$ and the similar argument as Case (I)–(III) to see that the contributions from first two terms in (4) to (4.3) are of $O_{t,n}(1)$. The contributions from last two terms in (4) are of $O_{t,n}(1)$ by Cauchy–Schwarz inequality, the assumption (1.8), Corollary 2.3 and Lemma 4.1. For the third term in (4), it follows from Cauchy–Schwarz inequality that
$\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{2n}x^{2nk(\sigma_{x}-1/2)}\bigg(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma\bigg)^{2n}$
$\leq\bigg(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{4n}x^{4nk(\sigma_{x}-1/2)}\bigg)^{1/2}$
$\times\bigg(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma\bigg)^{4n}\bigg)^{1/2},$ (4.9)
where $1\leq k\leq 7$. By Hölder’s inequality with the exponents 4n/(4n1)4n/(4n-1) and $4n$,
$\begin{split}&\bigg(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma\bigg)^{4n}\\ \leq&\,(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\mathrm{d}\sigma\bigg)^{4n-1}\bigg(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|^{4n}\mathrm{d}\sigma\bigg)\\ =&\,\frac{1}{(k\log x)^{4n-1}}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|^{4n}\mathrm{d}\sigma.\end{split}$
Note that
$\begin{split}&\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|^{4n}\\ &\ll_{n}\bigg(\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}+\sum_{p^{k}\leq x^{3}}\frac{\overline{C_{F}(p^{k})}\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma-\mathrm{i}t)}}\bigg)^{4n}\\ &\quad+\bigg(\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}-\sum_{p^{k}\leq x^{3}}\frac{\overline{C_{F}(p^{k})}\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma-\mathrm{i}t)}}\bigg)^{4n}.\end{split}$
By using the same argument as in proving (4.2), for $1\leq k\leq 7$, we get
$\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|^{4n}&\ll_{t,n}(\log x)^{8n}.\end{split}$
Thus,
$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma\bigg)^{4n}\\ &\ll_{t,n}\,\frac{1}{(k\log x)^{4n-1}}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}(\log x)^{8n}\mathrm{d}\sigma\ll_{t,n}(\log x)^{4n}.\end{split}$
Inserting this bound to (4) and using Lemma 4.1, we have the third term in (4) is of $O_{t,n}(1)$. This completes the proof of the lemma. ∎
Proof of Theorem 1.1. By the binomial theorem, we have
$\begin{split}S_{F}(t)^{n}=\mathcal{M}_{F}(t)^{n}+O_{n}\bigg(\sum_{\ell=1}^{n}|\mathcal{M}_{F}(t)|^{n-\ell}|\mathcal{R}_{F}(t)|^{\ell}\bigg),\end{split}$ (4.10)
where $\mathcal{M}_{F}(t)$ and $\mathcal{R}_{F}(t)$ are defined in (4.1). For $1\leq\ell\leq n$, we apply the generalized Hölder’s inequality, Corollary 2.3 and Lemma 4.3 to get
$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{M}_{F}(t)|^{n-\ell}|\mathcal{R}_{F}(t)|^{\ell}\\ \ll&\bigg(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg)^{\frac{1}{2}}\bigg(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{M}_{F}(t)|^{2n}\bigg)^{\frac{n-\ell}{2n}}\\ &\times\bigg(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{R}_{F}(t)|^{2n}\bigg)^{\frac{\ell}{2n}}\\ \ll&{}_{t,n}\,(\log\log T)^{(n-\ell)/2}\ll_{t,n}(\log\log T)^{(n-1)/2}.\end{split}$
Here we have used the positivity of $\mathcal{N}_{F}$ and the fact that $h_{T,M}(\mu_{F})\geq 0$ for all $\mathrm{GL}_{3}$ Hecke–Maass cusp forms $F$. Then (1.11) follows from (4.2), (4.10) and the above estimate.
For the second assertion of Theorem 1.1, note that the $n$-th moment of $\mu_{TM}$ is
$\int_{\mathbb{R}}\xi^{n}\mathrm{d}\mu_{TM}(\xi)=\bigg(\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg(\frac{S_{F}(t)}{\sqrt{\log\log T}}\bigg)^{n}\bigg)\bigg/\bigg(\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg).$
Using (1.11) and Corollary 2.3, we obtain that for all $n\geq 0$,
$\lim_{T\rightarrow\infty}\int_{\mathbb{R}}\xi^{n}\mathrm{d}\mu_{TM}(\xi)=C_{n}=\sqrt{\pi}\int_{\mathbb{R}}\xi^{n}\exp(-\pi^{2}\xi^{2})\mathrm{d}\xi.$
Then by the theory of moments in probability theory (see, for example, [3, Theorem 30.2]), we complete the proof of Theorem 1.1.
5. Conclusion
In conclusion, we establish an unconditional asymptotic formula for the moments of $S_{F}(t)$, where $F$ is a Hecke–Maass cusp form for $\mathrm{SL}_{3}(\mathbb{Z})$ with Langlands parameter in generic position. Our main result, Theorem 1.1, extends the conditional result of Liu and Liu [15], which assumed the GRH, to an unconditional setting. This achievement is made possible by applying a novel weighted zero-density estimate for $L(s,F)$ in the spectral aspect, recently established in [20]. Moreover, we demonstrate that the normalized values $S_{F}(t)/\sqrt{\log\log T}$ converge in distribution to a Gaussian random variable as $T\to\infty$. A central goal of our ongoing work is to extend this unconditional framework to $\mathrm{SL}_{n}(\mathbb{Z})$ for all $n\geq 3$, thereby providing an unconditional counterpart to the conditional theorem of Chen et al. [9].
Acknowledgements Q. Sun was partially supported by the National Natural Science Foundation of China (Grant Nos. 12471005 and 12031008) and the Natural Science Foundation of Shandong Province (Grant No. ZR2023MA003). The authors are very grateful to the referees for their valuable suggestions.